OpenStudy (anonymous):

What is -5 - 6 for g=-2?? Evaluated

OpenStudy (smokeysthename):

the question does not make sense. what are you trying to find?

OpenStudy (doc.brown):

Is there any more to that question?

OpenStudy (anonymous):

Well no. Im trying to evaluate it... I'm lost too,

OpenStudy (smokeysthename):

if its giving you g=-2 then there has to be more to the equation. you might be missing something.

OpenStudy (doc.brown):

(-5)+(-6) = (-11) If g = (-2) then g should be placed into an equation somewhere. Such as: -5g-6g =(-5)(-2)+(-6)(-2) =10+12 =22

OpenStudy (anonymous):

Well here are my Nswer choices..... 4, or 16 or 10, or -13

OpenStudy (smokeysthename):

go back and check to make sure you're not missing anything.

OpenStudy (doc.brown):

Is the question \[-5-g\]

OpenStudy (smokeysthename):

@doc.brown it can't be -5-g because -5-(-2) is not one of the answer choices

OpenStudy (anonymous):

This is exactly how it written out.... Evaluate -5g - for g=-2 that's what it says. No further information.....

OpenStudy (smokeysthename):

makes more since now. -5(-2) = 10

OpenStudy (smokeysthename):

sense*

OpenStudy (anonymous):

Oops forgot a six in between - and for

OpenStudy (smokeysthename):

okay -5(-2)-6= 4

OpenStudy (smokeysthename):

you have to plug in -2 for g.

OpenStudy (smokeysthename):

if you look at what you originally posted, it says -5-6 that's why we were confused

OpenStudy (anonymous):

So it's like this -5g - 6 for g= -2

OpenStudy (doc.brown):

\[-5g-6\]\[-5(-2)-6\]\[10-6\]\[=4\]

OpenStudy (anonymous):

Oh okay get it now!!! Thanks so much guys you guys are life savers!!!!

OpenStudy (smokeysthename):

No problem! if you have any other questions just post them :)